# Ditto: Motion-Space Diffusion for Controllable Realtime Talking Head Synthesis

### v2

@@ 4.3 Quantitative Evaluation
The quantitative results are shown in Table 1. Our method achieves the best results among competing methods on all evaluated metrics. The compact motion representation facilitates the learning of the diffusion model and makes the lipsync score close to the real video, which is the most important criterion for evaluating talking head tasks. How to maintain identity consistency, measured by the CSIM score, is a long-standing challenge for one-shot methods. Our method achieves higher CSIM scores than other methods, attributed to the disentanglement of appearance and motion. Our method also effectively improves the frame quality and inter-frame continuity of the generated video obtaining the lowest FID and FVD among all methods. Table 1 also reports the results of reducing the number of denoising steps. The results of 10-step and 50-step denoising exhibit minimal disparity in end-to-end evaluation, substantiating the efficacy of our inference optimization. Analysis of the RTF reveals that diffusion-based methods exhibit inference speeds 30-50 times slower than realtime. In contrast, our proposed method achieves inference capabilities that surpass realtime performance, and even outpace the non-diffusion-based lip-editing method MuseTalk.
@@ 4.4 Qualitative Evaluation
Figure 4 shows a qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generate videos using the same audio input with each method and selected frames at the same location for comparison. Our method shows a notable edge in the consistency of generated details. For example, the shape, number and texture of teeth generated by other methods may change between frames, while our generation results show consistency between consecutive frames. The comparison methods also have problems such as generating exaggerated expressions, blur, and inaccurate lip movements. These problems are unacceptable concerns to customers in real application scenarios. In addition, the generation results of Hallo and Hallo2 for stylized portraits may deteriorate gradually over time, such as excessive facial wrinkles, abnormal shadows, and background artifacts. In contrast, our method demonstrates consistent performance across diverse portrait styles.
@@ 4.5 Ablation Study
Adaptive Loss Weights. Without adaptive loss weights (w/o Ada-w), the problem of unbalanced convergence between pose and expression becomes severe, hindering the model’s ability to learn the correspondence between audio and facial motion. We observe that the generated videos exhibit minimal facial motions, resulting in poor Sync and FVD scores, while CSIM and FID metrics are less affected.
4.6 Inference Performance
@@ 5 Conclusion
We propose a diffusion-based framework specifically designed for one-shot talking head synthesis. We use a Diffusion Transformer to generate motion sequences in an identity-independent space and synthesize videos through a dedicated face renderer. It can produce realistic videos from a single portrait image and support fine-grained motion control, promising for realtime interactive applications.

### v3

@@ 4.2 Baseline and Metrics
Table 3: User study results.
Method Visual quality Lipsync Naturalness
EchoMimic 54.0% 38.0% 58.7%
Hallo 21.3% 36.0% 33.3%
Hallo2 40.7% 45.3% 59.3%
Ours 84.0% 80.7% 48.7%
Figure 5: The qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generate videos using the same audio input with each method and select frames at the same location for comparison. Blue arrows indicate the locations of artifacts, including inaccurate lip movements, disordered teeth, blur, etc.
@@ 4.3 Quantitative Evaluation
The quantitative results on Talk9 and HDTF100 datasets are shown in Table 1 and Table 2, respectively. Our method achieves the best results among competing methods on all evaluated metrics. The motion representation facilitates the learning of the diffusion model and makes the lipsync score close to the real video, which is the most important criterion for evaluating talking head tasks. How to maintain identity consistency, measured by the CSIM score, is a long-standing challenge for one-shot methods. Our method achieves higher CSIM scores than other methods, attributed to the disentanglement of appearance and motion. Our method also effectively improves the frame quality and inter-frame continuity of the generated video obtaining the lowest FID and FVD among all methods. Table 1 also reports the results of reducing the number of denoising steps. The results of 10-step and 50-step denoising exhibit minimal disparity in end-to-end evaluation, substantiating the efficacy of our inference optimization. Analysis of the RTF reveals that diffusion-based methods exhibit inference speeds 30-50 times slower than realtime. In contrast, our proposed method achieves inference capabilities that surpass realtime performance, and even outpace the non-diffusion-based lip-editing method MuseTalk.
User Study.
To estimate the quality of our method and SOTAs from human perspectives, we conduct a blind user study with 10 participants. Specifically, we randomly select 5 image-audio pairs to generate 5 driving clips for each tested method, resulting in a total of 20 clips. Each participant is presented with two videos generated by different methods for the same set of inputs, and asked to choose the better one in terms of visual quality, lip synchronization, and naturalness of facial expressions and head movements. This process is repeated $C_{4}^{2}$ times. The results are summarized in Table 3, where our method outperforms other methods in visual quality and lip synchronization. We also notice that Ditto performs slightly weaker in the measure of “naturalness of facial expressions and head movement”. We believe this is due to the motion representation attenuating the learning of some high-frequency motion information, which still requires our continuous optimization.
@@ 4.4 Qualitative Evaluation
The qualitative results are presented in Figures 4 and Figure 5. Figure 5 shows a qualitative comparison of our approach on two characters with different styles, poses, and hairstyles. For each character, we generate videos using the same audio input with each method and select frames at the same location for comparison. Our method shows advantages in the consistency of generated details. For example, the shape and texture of teeth generated by baselines change between frames. The baselines also have problems such as generating exaggerated expressions, blur, and inaccurate lip movements. In addition, the generation results of Hallo and Hallo2 for stylized portraits may deteriorate gradually over time, such as excessive facial wrinkles, abnormal shadows, and background artifacts. In contrast, our method shows stable performance across diverse portrait styles.
@@
Figure 6: The gaze adjustment decouples gaze direction from the head pose, enabling the avatar’s eyes to focus on the camera, producing more natural and engaging eye contact.
@@ 4.5 Ablation Study
Adaptive Loss Weights. Without adaptive loss weights (w/o Ada-w), the problem of unbalanced convergence between pose and expression becomes severe, hindering the model to learn the correspondence between audio and facial motion. We observe that the generated videos exhibit minimal facial motions, resulting in poor Sync and FVD scores, while CSIM and FID metrics are less affected.
Gaze Adjustment. As shown in Figure 6, without gaze adjustment, the gaze direction is rigidly bound to the head pose. The gaze adjustment decouples gaze direction from the head pose, enabling the avatar’s eyes to focus on the camera, producing more natural and engaging eye contact.
@@ 5 Conclusion
We propose Ditto, a diffusion-based talking head framework that enables fine-grained controls and real-time inference. We devise a Conditional DiT and corresponding training strategy to address the issues in generating motion representations. We jointly optimize the holistic framework to enable streaming processing, real-time inference, and low first-frame delay. For the advancement of the community, we make our source code available for open-source use.